The 2017 balance sheet of Dream, Inc., showed current assets of $4,515 and current liabilities of $3,005. The 2018 balance sheet showed current assets of $2,990 and current liabilities of $1,590.

What was the company’s 2018 change in net working capital, or NWC? (Do not round intermediate calculations. A negative answer should be indicated by a minus sign.)
Answer #1

2017 NWC = Current assets – current liabilities

=4515-3005

=1510

2018 NWC = 2990-1590 = 1400

2018 Change in NWC = 2018 NWC – 2017 NWC

= 1400- 1510

= -$110